OpenStudy (anonymous):

Which of the following is the solution set of 18 - 3n + 2 = n + 20 - 4n? Ø {0} all reals

OpenStudy (anonymous):

18 - 3n + 2 = n + 20 - 4n 20-3n=20-3n Now, substitute any value for n and see u get a True statement or False statement.

OpenStudy (anonymous):

Simplifying 18 + -3n + 2 = n + 20 + -4n Reorder the terms: 18 + 2 + -3n = n + 20 + -4n Combine like terms: 18 + 2 = 20 20 + -3n = n + 20 + -4n Reorder the terms: 20 + -3n = 20 + n + -4n Combine like terms: n + -4n = -3n 20 + -3n = 20 + -3n Add '-20' to each side of the equation. 20 + -20 + -3n = 20 + -20 + -3n Combine like terms: 20 + -20 = 0 0 + -3n = 20 + -20 + -3n -3n = 20 + -20 + -3n Combine like terms: 20 + -20 = 0 -3n = 0 + -3n -3n = -3n Add '3n' to each side of the equation. -3n + 3n = -3n + 3n Combine like terms: -3n + 3n = 0 0 = -3n + 3n Combine like terms: -3n + 3n = 0 0 = 0 Solving 0 = 0 Couldn't find a variable to solve for. This equation is an identity, all real numbers are solutions.